OpenStudy (anonymous):

Match each system of linear equations with the quadrant in which the solution lies. Answer two y plus sixteen equals negative three x...y minus one forth x equals five Read Answer Items for Question 14 Y minus fifty equals ten x...y minus x equals two Read Answer Items for Question 14 Five y plus one equals two x...negative two y equals x minus fourteen Read Answer Items for Question 14 Y plus three x equals eight...seventeen plus y equals five x Read Answer Items for Question 14 Answer A. Two B. Four C. Three D. One Match each system of linear equations with the quadrant in which the solution lies. Answer two y plus sixteen equals negative three x...y minus one forth x equals five Read Answer Items for Question 14 Y minus fifty equals ten x...y minus x equals two Read Answer Items for Question 14 Five y plus one equals two x...negative two y equals x minus fourteen Read Answer Items for Question 14 Y plus three x equals eight...seventeen plus y equals five x Read Answer Items for Question 14 Answer A. Two B. Four C. Three D. One @Mathematics

OpenStudy (anonymous):

please help me!

OpenStudy (anonymous):

the answers are the different quadrants if that helps any

OpenStudy (turingtest):

y+16=-3x y-x/4=5--->y=x/4+5 x/4+5+16=-3x 13x/4=-21 x=-84/13 y=-21/13+5=44/13 this is in the second quadrant

OpenStudy (turingtest):

because we have a positive y and negative x...the others work the same way, but I may not have time to finish, I'm leaving soon.

OpenStudy (turingtest):

in general if the solution to a system has y>0 and x>0 we are in quadrant 1 y>0 and x<0 we are in 2 y<0 and x<0 we are in 3 y<0 and x>0 we are in 4

OpenStudy (turingtest):

next one: y-50=10x y-x=2--->y=x+2 x+2-50=10x 9x=-48 x=-48/6<0 y=-48/6+12/6=-6<0 so since y<0 and x<0 we are in quadrant 3

OpenStudy (anonymous):

i just dont understand how to take an equation and put it to a graph. Do you get what im saying?

OpenStudy (turingtest):

you cannot graph y=x?

OpenStudy (turingtest):

can you graph any equation? or are you completely lost in that department?

OpenStudy (anonymous):

this is my first time ever seeing this sort of stuff

OpenStudy (anonymous):

so yea sorta completely lost see i am previewing a pretest that leads into the lessons for this stuff.

OpenStudy (turingtest):

well if you know practically nothing of graphing I don't know why you are solving solution sets, that is more advanced than just graphing y=x I will try to give the idea briefly: to graph a linear function we need to find two points and draw a straight line. to find our points use the equation and try various values of x: let's use the example y=2x+1 if x=0 then y=2(0)+1=1 so the point (0,1) is on our graph if x=1 y=2(1)+1=5 so (1,5) is another point. do you know how to plot these points on a graph?

OpenStudy (anonymous):

oo ok yea i know how to plot. and thanks i understand it now

OpenStudy (turingtest):

just because I had almost finished the drawing anyway, here it is (for the example I created)